Given f(x) = x^2 + x and h(x) = x - 2, find ( f O h)(x)
Answer by HyperBrain(694)

Given f(x) = x^2 + x and h(x) = x - 2
: replace all x's in f(x) with x-2
f(g(x))=(x-2)2+x-2
f(g(x))=(x2-4x+4)+x-2
f(g(x))=x2-3x+2
Power up,
HyperBrain!
bored?
yeah!